OpenStudy (anonymous):

Mallorie has $3 in her wallet. If this is 10% of her monthly allowance, what is her monthly allowance? This is talking about Percent Proportion. How do you find The parts, whole, and percents in a equation?

OpenStudy (shamil98):

\[\huge \frac{ 3 }{ .10} = 30\] 1/10 = 0.10

OpenStudy (btaylor):

The part she has is $3. The percent is 10%. So, \[\frac{part}{whole} \times 100 = percent\]

OpenStudy (btaylor):

Put in what you know (The part and percent) and solve for what you don't.

OpenStudy (anonymous):